Help with this and I give thumbs up please
his Question. T p Oldhat Financial starts its first day of operations with $11 million in capital. A total of $125 million in checkable deposits are received. The bank makes a $25 million commercial loan and another $60 million in mortgages, with the following terms: 200 standard 30-year fixed-rate mortgages with a nominal annual rate of 5 25%, each for S300 000 Assume that required reserves are 896. Complete the banks balance sheet provided below. (Round your responses to the nearest whole number) Assets million Checkable deposits millionBank capital million S□ million Excess reserves Loans S million The leverage ratio is 96, and the bank is ! The risk-weighted assets after Oldhats first day are S□ millon Round your esponse to the nearest whole number) The risk-weighted capital ratio after Oldhats first day is □% (Round your response to two decimal places) (Round your response to two decimal places)
Answer #1

Answer.)

Deposits received = $125 million

Reserve requirement=8%

So Required reserves=0.08*125=$10 million

Loans=commercial loans+mortgage loans=25+60=$85 million

Excess Reserves=(deposit+capital-loans)-required reserves=(125+11-85)-10=$41 million

Checkable Deposits=$125 miilion

Bank capital=$11 million

Leverage Ratio=bank's capital/total liabilities=9/134=6.7% and the bank is well capitalized

Risk weighted assets of first day:

Reserve weightage is 0

Commercial loans weightage = 100%

Risk Weighted assets=$5 million

mortgage loans weightage=50%

Risk Weighted assets=$30 million

Total RW Assets=$35 million

So, Risk Weighted capital ratio=capital/RW assets=11/35=31.43%